Write the first five terms of the arithmetic sequence. Find the common difference and the nth term of the sequence as a function of n.
a 1=2; a k+1=a k-4
I'm a very confused about this question. I understand the Fibonnaci sequence though. Do I solve this the same way?

Answer by richard1234(7193)

Given a_1 = 2, a_(k+1) = a_k - 4, we have
a_1 = 2
a_2 = a_1 - 4 = 2 - 4 = -2
a_3 = a_2 - 4 = -2 - 4 = -6.
Therefore the common difference is -4, and we can establish:
a_n = 6 - 4n (based on the original conditions)
Also, the Fibonacci sequence is not an arithmetic sequence, as the difference between successive terms increases.